Rogelio’s portfolio is worth 69,400 dollars and has three stocks. It has 23,100 dollars of stock A, which has an expected return of 6.86 percent; it has 1,400 shares of stock B, which has a share price of 5 dollars and an expected return of 7.21 percent; and it has some stock C, which has an expected return of 13.58 percent. What is the expected return of Rogelio’s portfolio? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

We can calculate the expected return of a portfolio using the following formula:-

= Wa*Er(a)+Wb*Er(b)+Wc*Er(c)

Wa = weight of a = 23,100/69400 = 0.3328

Er(a) = Expected return of a =6.86%

Wb = weight of b =1400*5 = 7000/69400= 0.10

Er(b) = Expected return of b = 7.21%

Wc = weight of c = 1-0.3328-0.10 = 0.5672

Er(c) = Expected return of c = 13.58%

= 0.3328*6.86+0.10*7.21+0.5672*13.58

= 2.283+0.721+7.702

Expected return of portfolio = 10.706%